$SPAD/src/input parsurf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
-- test for ParametricSurface domain
)clear all
f1(u:DFLOAT,v:DFLOAT):DFLOAT == u*cos(v)
f2(u:DFLOAT,v:DFLOAT):DFLOAT == u*sin(v)
f3(u:DFLOAT,v:DFLOAT):DFLOAT == cos(u+v)
draw(surface(f1,f2,f3),0..2*%pi,-%pi..%pi)
f4(u:DFLOAT,v:DFLOAT):DFLOAT == v*cos(u+v)
draw(surface(f1,f2,f4),0..2*%pi,-%pi..%pi)

———-

References^{[1] nothing}

4